If f(x)=2(3^x-2),what is the value of f(2)?
Answer by checkley71(8403)

YOU ARE CORRECT IF F(X)=2(3^[X-2])
MY ANSWER IS CORRECT IF F(X)=2(3^X)-2
F(X)=2(3^X-2)
F(2)=2(3^2-2)
F(2)=2(9-2)
F(2)=2(7)
F(2)=14 ANSWER.